Rock A is dropped straight down. After 2 seconds, Rock B is dropped straight down as well. As both rocks are falling, will the separation distance between Rock A and B increase, decrease or stay the same?
1
Expert's answer
2012-11-19T07:40:05-0500
1 rock's equation: h1(t)=h0-gt^2/2
2 rock's equation: h2(t)=h0-g(t-2)^2/2
d(t)=h1(t)-h2(t)=g/2(t^2-t^2+4t-4)=2g(t-1)
So, it is linearly increasing